OpenStudy (anonymous):

Graph a hyperbola with foci at (0, ±10) and vertices at (0, ±8) using graphing technology. PLEASE HELP ME! I CANT GO TO SLEEP UNTIL I FINISH THIS :(

OpenStudy (anonymous):

my graphing technology is geogbra by the way

OpenStudy (anonymous):

so the actual question is to come up with the equation of the hyperbola with those as the given info so you can graph it in geogebra, right?

OpenStudy (anonymous):

since the foci are on a vertical line, your hyperbola will be the up/down type... the standard equation for this type of hyperbola is: \(\large \frac{(y-k)^2}{a^2}-\frac{(x-h)^2}{b^2}=1 \) where (h, k) is the center of your hyperbola, and your vertices are a units above and below this center. b is obtained by the rlation: \(\large c^2=a^2+b^2 \) where c is the distance from the center to each of your foci.

OpenStudy (anonymous):

okay. Im still confused though. I dont have any of the things I need to graph it. right?

OpenStudy (anonymous):

Can anyone help me!?

OpenStudy (anonymous):

Try reading through the lesson and following the directions or call your instructor if you still need help

OpenStudy (anonymous):

I cant call him yet though. its too early and Im stuck. Theres no examples in the lesson like this one

OpenStudy (anonymous):

on page 2, example #2. You are given the verticies, so you can figure out the length of the transverse axis